# Extended formulations for induced tree and path polytopes of chordal graphs

### v2

@@ Abstract
In this article, we give two extended space formulations, respectively, for the induced tree and path polytopes of chordal graphs with vertex and edge variables. These formulations are obtained by proving that the induced tree and path extended incidence vectors of chordal graphs form Hilbert basis. This also shows that both polytopes have the integer decomposition property in chordal graphs.
Whereas the formulation for the induced tree polytope is easily seen to have a compact size, the system we provide for the induced path polytope has an exponential number of inequalities. We show which of these inequalities define facets and exhibit a superset of the facet-defining ones that can be enumerated in polynomial time. We show that for some graphs, the latter superset contains redundant inequalities. As corollaries, we obtain that the problems of finding an induced tree or path maximizing a linear function over the edges and vertices are solvable in polynomial time for the class of chordal graphs.
@@
All the graphs in this article are simple and connected unless specified. Given a graph $G=(V,E)$, we denote its complement by $\overline{G}=(V,\overline{E})$, where $\overline{E}=\{e\in\binom{V}{2}:e\notin E\}$. We respectively denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$. Two vertices $u$ and $v$ are adjacent if $uv\in E(G)$. Given a subset of vertices $W\subseteq V$, let $E(W)$ denote the set of edges of $G$ having both endpoints in $W$ and $\delta(W)$ the set of all edges having exactly one endpoint in $W$. When $W$ is a singleton $\{w\}$, we will simply write $\delta(w)$. We say that the edges in $\delta(w)$ are incident to $w$, and two edges sharing an endpoint are said to be adjacent. Given two subsets of vertices $W_{1}$, $W_{2}$, we denote by $\delta(W_{1},W_{2})$ the set of edges having one endpoint in both $W_{1}$ and $W_{2}$. For $F\subseteq E$, let $V(F)$ denote the set of vertices incident to any edge of $F$. Given $W\subseteq V$, the graph $G[W]=(W,E(W))$ is the subgraph of $G$ induced by $W$. When $H$ is an induced subgraph of $G$, we say that $G$ contains $H$. Given a vertex $u\in V(G)$, we denote by $N_{G}(u)=\{w\in V(G):uw\in E(G)\}$ its neighborhood, and by $N_{G}[u]=N_{G}(u)\cup\{u\}$ its closed neighborhood, when clear from context we simply write $N(u)$ and $N[u]$. Two vertices $u$ and $v$ are true twins if $N[u]=N[v]$ and false twins if $N(u)=N(v)$. A vertex $w$ is complete to a subset of vertices $W$ if it is adjacent to each vertex of $W$ and does not belong to $W$. A vertex $u$ complete to $V\setminus\{u\}$ is said to be universal. Given a set of vertices $W\subseteq V$, we denote by $C_{G}(W)$ the vertices complete to $W$ in $G$; we simply write $C(W)$ when clear from context. A vertex is simplicial if its neighborhood induces a clique.
@@
An induced path (resp. hole) is a graph induced by a sequence of vertices $(v_{1},\dots,v_{p})$ whose edge set is $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}$ (resp. $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}\cup\{v_{1}v_{p}\}$ ). Note that sets of vertices inducing cliques with three vertices are considered holes. An induced tree is a set of vertices inducing an acyclic graph. A subset of vertices induces a path (resp. hole) if its elements can be ordered into a sequence inducing a path (resp. hole). The length of a path or hole is its number of edges. A path or hole is called long if its length is at least 4. The parity of a path or hole is the parity of its length. Given a path induced by the sequence $(w_{1},\dots,w_{k})$, we call $w_{1}$, $w_{k}$ its extremities and the set of vertices $\{w_{2},\dots,w_{k-1}\}$ its interior. A tree is a graph with no hole.
The contraction of an edge $uv$ in $G$ yields a new graph $G/uv$ built from $G$ by deleting $u$, $v$, adding a new vertex $w$, and adding the edges $wz$ for all $z\in N_{G}(u)\cup N_{G}(v)$; parallel edges and loops are removed. For $F\subseteq E$, we denote by $G/F$ the graph obtained from $G$ by contracting all the edges in $F$.
@@
A graph is $k$-chordal if its holes have maximum size $k$; 3-chordal graphs are called chordal. Equivalently, chordal graphs are those for which an order $(v_{1},\dots,v_{k})$ of its vertices exists such that $v_{i}$ is simplicial in $G_{i}=G[\{v_{i},\dots,v_{k}\}]$, such an order is called a perfect elimination order. Moreover, the maximal cliques of $G$ are among $N_{G_{i}}[v_{i}]$, which shows that they come in a linear number.
@@
The size of a linear system is its number of rows. Let $\mathcal{P}=\{x:Ax\leq b\}\subseteq\mathds{R}^{n}$ be a polyhedron, that is, the intersection of finitely many half-spaces. The dimension of $\mathcal{P}\subseteq\mathds{R}^{n}$, denoted by $\textrm{dim}~\mathcal{P}$, is the maximum number of affinely independent points in $\mathcal{P}$ minus one. If $a\in\mathds{R}^{n}\setminus\{0\}$, $\alpha\in\mathds{R}$, then the inequality $a^{\top}x\leq\alpha$ is said to be valid for $\mathcal{P}$ if $\mathcal{P}\subseteq\{x\in\mathds{R}^{n}:a^{\top}x\leq\alpha\}.$ We say that a valid inequality $a^{\top}x\leq\alpha$ defines a facet of $\mathcal{P}$ if $\textrm{dim}(\mathcal{P}\cap\{x\in\mathds{R}^{n}:a^{\top}x=\alpha\})$ is equal to $\textrm{dim}~\mathcal{P}$ minus one. The polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is full dimensional when dim $\mathcal{P}$ is equal to $n$. In that case, a linear system $Ax\leq b$ that defines $\mathcal{P}$ is minimal if each inequality of the system defines a facet of $\mathcal{P}$. Moreover, these facet-defining inequalities are unique up to scalar multiplication. A face of $\mathcal{P}$ is a polyhedron contained in $\mathcal{P}$ obtained by setting to equality some inequalities in $Ax\leq b$. Equivalently, a face is the intersection of the supporting hyperplane of a valid inequality and $\mathcal{P}$. Extreme points of a polyhedron are its faces of dimension zero. Each extreme point satisfies $\textrm{dim}~\mathcal{P}$ linearly independent valid inequalities for $\mathcal{P}$ with equality. A polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is said to be integer if all its non-empty faces contain an integer point, it is binary if it is bounded and all its extreme points are binary. A polytope is a bounded polyhedron, equivalently, it can be defined as the convex hull of a finite number of points $P$ and is denoted by $\textrm{conv}(P)$. A cone is a polyhedron described by a linear system with a right hand side equal to 0; it is pointed if it contains no affine space. The elements of a cone can be described by conic combinations of finite sets of elements called generators. The minimal generating sets of pointed cones are unique up to positive scalar multiplication; we denote the element of such a family by its generators they are in one to one correspondence with its face of dimension 1. A polyhedron $\mathcal{P}$ is said to have the integer decomposition property (IDP) if every integer point in the dilation $k\mathcal{P}$ for $k\in\mathds{Z}_{+}$ is obtained as the sum of $k$ elements of $\mathcal{P}$. A set of points $X=\{\chi_{1},\dots,\chi_{n}\}$ forms a Hilbert basis if every integer point in the conic hull of $X$ is expressible as an integer nonnegative combination of elements of $X$. The Hilbert basis and integer decomposition property are closely related. If the extreme points of a polytope $\mathcal{P}$ form a Hilbert basis, then $\mathcal{P}$ has the IDP.
@@
In this work, the polyhedra we consider are convex hulls of points encoding structured induced subgraphs of an input graph $G=(V,E)$. Such an induced subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ (note that $E^{\prime}=E(V^{\prime})$) can be encoded by its natural incidence vector that is $\chi^{V^{\prime}}\in\{0,1\}^{V}$ whose components associated with elements of $V^{\prime}$ are equal to 1 and the others equal to 0, or by its extended incidence vector that is $\xi^{V^{\prime}}=(\chi^{V^{\prime}},\zeta^{E^{\prime}})\in\{0,1\}^{V\cup E}$ where the components of $\zeta^{E^{\prime}}$ associated with elements of $E$ are equal to 1 and the others equal to 0. Given a point $(x,y)\in\mathds{R}^{V\cup E}$ we denote by $(x,y)|_{G^{\prime}}$ its restriction to elements of $G^{\prime}$.
In this paper, given a graph $G$, we consider the following polyhedra:
  • its induced tree cone $\text{cone}\{\xi^{W}:\text{ $W$ induces a tree of $G$}\}$
  • its induced tree polytope $\text{conv}\{\xi^{W}:\text{ $W$ induces a tree of $G$}\}$
  • its induced path cone $\text{cone}\{\xi^{W}:\text{ $W$ induces a path of $G$}\}$
  • its induced path polytope $\text{conv}\{\xi^{W}:\text{ $W$ induces a path of $G$}\}$
@@
In this paper, we study the problems of finding maximum weight induced trees and paths where both the edges and the vertices have weight, and in particular, filling a gap in polyhedral literature associated with these problems.
@@
The problem of finding the size of the largest induced tree in a graph is a problem that has been approached through extremal combinatorics [ERDOS198661, MATOUSEK2007307, FOX2009494]. In particular, an important question is that of bounding this number according to several graph parameters, such as the number of vertices, edges, radius, independence number, maximum clique, and connectivity. In [ERDOS198661], a Ramsey-type bound is proposed on the minimum integer $n(k,t)$ such that every graph having at least $n(k,t)$ vertices contains either a clique of size $k$ or an induced tree of size $t$. Finally, they show that every connected graph with radius $r$ has an induced tree of size at least 2r12r-1 by exhibiting a large induced path. A related problem is the one of finding large induced forests [inducedforest] in which bounds for the largest induced forest are given with respect to maximum degree or stability number. Besides these studies, no polyhedral study of the induced tree polytope on particular classes has been proposed. The problem of computing the maximum size of an induced tree is known to be NP-hard [completePathTree] in general, but it is also the case in bipartite and triangle-free graphs with maximal degree 3 [MATOUSEK2007307]. Several integer linear programs have been proposed to find maximum weighted such subsets [ILPinducedtree], and extensive computational experiments have been made on general graphs.
@@
The problem of finding a longest induced path has applications in assessing the robustness of a network. It can be viewed as a variant of the graph diameter. In a graph where vertices represent communication devices that may fail, and edges represent possible transmission links, the length of the longest induced path indicates the worst-case number of intermediate devices required for a one-to-one communication between any two vertices $u$ and $v$ after the failure of a set of devices that do not disconnect $u$ from $v$. Another application is about error correcting codes [error]. The induced paths of a graph are induced trees, and the type of questions raised in the previous paragraph can be raised considering this structure. The authors of [FOX2009494] show that every connected triangle-free graph on $n$ vertices has an induced path of size at least $\sqrt{n}$, and their approach can be generalized to obtain lower bounds for graphs with no clique of given size $k$. In [esperet], the authors show that every 3-connected planar graph contains an induced path on $\text{log}(|V|)$ vertices. In [jfraymond], it is observed that a graph may contain a large path but no large induced path; for this reason, they characterize the sets of edges to be forbidden along a path of size $n$ to force the existence of an induced path of size $log(n)$. This result allows them to show that if a graph has a path of length $n$ and no $K_{t}$ as a topological minor, then it contains an induced path of order $(logn)^{\Omega(1/tlog^{2}t)}$.
Finding a maximum vertex weighted induced path is also NP-hard [completePathTree], but polynomially solvable in $k$-chordal graphs, interval filament, graphs decomposable by clique cutset or by splits whose prime subgraphs are polynomially solvable [GAVRIL2002203], improved in [HANSEN2009135] and on graphs of bounded clique width, asteroidal triple free graphs [inducedpathcliquewifth]. Besides these combinatorial algorithms, no polyhedral study of the corresponding polytope has been made. A related polyhedra has, however, been characterized in [goel2025halfspacerepresentationspathpolytopes]: the polytope whose extreme points are incidence vectors of paths between pairs of leaves in a tree: in trees, induced path and path are the same subgraphs, so the latter paper describes the induced path (between leaves) polytope of trees. Finally, extended space integer linear programs for this problem have been proposed [ILPinducedPath, ILPinducedPath2], leading to computational experiments.
@@
In this paper, we study both the induced tree polytope and the induced path polytope of chordal graphs in two separate sections. It turns out that beside being included one in the other, these two polytopes have similar properties. Each section starts by respectively showing that the induced path and tree extended incidence vectors of chordal graphs form a Hilbert basis of some cones, which we explicitly describe by linear systems of inequalities. We show that intersecting these cones with the correct hyperplane yields the induced tree and path polytope of chordal graphs in extended variable space. While the system we provide for the induced tree polytope is easily seen to be compact in chordal graphs, the one we provide for the induced path polytope has an exponential number of inequalities. Therefore, we characterize the ones that define facets and show that a superset can be enumerated in polynomial time. This implies that there exists a compact system describing the induced path polytope of chordal graphs. Finally, these results imply that the problems of finding a maximum vertex and edge weight induced tree or path are polynomially solvable in chordal graphs.

### v3

@@ Abstract
In this article, we give an extended space formulation for the induced tree polytope and another one for the induced path polytope of chordal graphs, both with vertex and edge variables. These formulations are obtained by proving that, in chordal graphs, the extreme points of those polytopes respectively form two Hilbert bases. Since these polytopes are binary, it also shows that they have integer decomposition property.
While the formulation for the induced tree polytope has a polynomial size, the formulation we propose for the induced path polytope involves an exponential number of inequalities. We identify which of these inequalities are facet‑defining and exhibit a polynomially enumerable superset containing all of them, although this superset may include redundant inequalities for some graphs.
As corollaries, we obtain that the problems of finding an induced tree or path maximizing a linear function over the edges and vertices are solvable in polynomial time for the class of chordal graphs.
@@
All the graphs in this article are simple. Given a graph $G=(V,E)$, we denote its complement by $\overline{G}=(V,\overline{E})$, where $\overline{E}=\{e\in\binom{V}{2}:e\notin E\}$. We denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$, respectively. Two vertices $u$ and $v$ are adjacent if $uv\in E(G)$. Given a subset of vertices $W\subseteq V$, we denote by $E(W)$ the set of edges of $G$ having both endpoints in $W$, and by $\delta(W)$ the set of all edges having exactly one endpoint in $W$. When $W$ is a singleton $\{w\}$, we simply write $\delta(w)$. We say that the edges in $\delta(w)$ are incident to $w$. Given two subsets of vertices $W_{1}$, $W_{2}$, we denote by $\delta(W_{1},W_{2})$ the set of edges having one endpoint in $W_{1}$ and the other in $W_{2}$. For $F\subseteq E$, we denote by $V(F)$ the set of vertices incident to any edge of $F$. Given $W\subseteq V$, the graph $G[W]=(W,E(W))$ is the subgraph of $G$ induced by $W$. When $H$ is an induced subgraph of $G$, we say that $G$ contains $H$. We denote by $G\setminus v$ (resp. $G\setminus uv$¨) the obtained from $G$ by removing the vertex $v$ (resp. edge $uv$). Given a vertex $u\in V(G)$, we denote by $N_{G}(u)=\{w\in V(G):uw\in E(G)\}$ its neighborhood, and by $N_{G}[u]=N_{G}(u)\cup\{u\}$ its closed neighborhood, when clear from context we simply write $N(u)$ and $N[u]$. Two vertices $u$ and $v$ are true twins if $N[u]=N[v]$ and false twins if $N(u)=N(v)$. A vertex $w$ is complete to a subset of vertices $W$ if it is adjacent to each vertex of $W$ and does not belong to $W$. A vertex $u$ complete to $V\setminus\{u\}$ is said to be universal. Given a set of vertices $W\subseteq V$, we denote by $C_{G}(W)$ the vertices complete to $W$ in $G$; we simply write $C(W)$ when the graph is clear from context. A clique is a set of pairwise adjacent vertices. A stable set is a set of pairwise nonadjacent vertices. A graph whose vertex set can be partitioned into two stable sets is called bipartite. A vertex is simplicial if its neighborhood induces a clique.
A path (resp. hole) is a graph induced by a sequence of vertices $(v_{1},\dots,v_{p})$ whose edge set is $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}$ (resp. $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}\cup\{v_{1}v_{p}\}$ and $p\geq 4$). A cycle is a set of edges inducing a connected graph with vertices of degree equal to 2. A tree is an acyclic graph.
@@
A subset of vertices induces a path (resp. hole) if its elements can be ordered into a sequence inducing a path (resp. hole). By abuse of definition, we consider that a set of vertices inducing a path is an induced path, similarly, an induced tree is a set of vertices inducing a tree. The length of a path or hole is its number of edges. The parity of a path or hole is the parity of its length. Given a path induced by the sequence $(w_{1},\dots,w_{k})$, we call $w_{1}$, $w_{k}$ its extremities and the set of vertices $\{w_{2},\dots,w_{k-1}\}$ its interior moreover, we say that $P$ ends in $w_{1}$ and $w_{k}$. The contraction of an edge $uv$ in $G$ yields a new graph $G/uv$ built from $G$ by deleting $u$, $v$, adding a new vertex $w$, and adding the edges $wz$ for all $z\in N_{G}(u)\cup N_{G}(v)$; parallel edges and loops are removed. For $F\subseteq E$, we denote by $G/F$ the graph obtained from $G$ by contracting all the edges in $F$.
@@
A graph is $k$-chordal if its holes have maximum size $k$; graphs with no holes are called chordal. Equivalently, a graph is chordal if there exists an order $(v_{1},\dots,v_{k})$ of its vertices such that $v_{i}$ is simplicial in $G_{i}=G[\{v_{i},\dots,v_{k}\}]$; such an order is called a perfect elimination order. Moreover, the maximal cliques of $G$ are among $\{N_{G_{i}}[v_{i}]:i\in\{1,\dots,k\}\}$, which shows that their number is linear.
@@
Let $\mathcal{P}=\{x:Ax\leq b\}\subseteq\mathds{R}^{n}$ be a polyhedron, that is, the intersection of finitely many half-spaces. The dimension of $\mathcal{P}\subseteq\mathds{R}^{n}$, denoted by $\textrm{dim}~\mathcal{P}$, is the maximum number of affinely independent points in $\mathcal{P}$ minus one. If $a\in\mathds{R}^{n}\setminus\{0\}$, $\alpha\in\mathds{R}$, then the inequality $a^{\top}x\leq\alpha$ is said to be valid for $\mathcal{P}$ if $\mathcal{P}\subseteq\{x\in\mathds{R}^{n}:a^{\top}x\leq\alpha\}.$ We say that a valid inequality $a^{\top}x\leq\alpha$ defines a facet of $\mathcal{P}$ if $\textrm{dim}(\mathcal{P}\cap\{x\in\mathds{R}^{n}:a^{\top}x=\alpha\})$ is equal to $\textrm{dim}~\mathcal{P}$ minus one. The polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is full dimensional when dim $\mathcal{P}$ is equal to $n$. In that case, a linear system $Ax\leq b$ that defines $\mathcal{P}$ is minimal if each inequality of the system defines a facet of $\mathcal{P}$. Moreover, these facet-defining inequalities are unique up to scalar multiplication. The size of a linear system of inequalities is its number of rows. A face of $\mathcal{P}$ is a polyhedron contained in $\mathcal{P}$ obtained by setting to equality some inequalities in $Ax\leq b$. Equivalently, a face is the intersection of the supporting hyperplane of a valid inequality and $\mathcal{P}$. Extreme points of a polyhedron are its faces of dimension zero. Each extreme point satisfies $\textrm{dim}~\mathcal{P}$ linearly independent valid inequalities for $\mathcal{P}$ with equality. A polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is said to be integer if all its non-empty faces contain an integer point. A polytope is a bounded polyhedron, equivalently, it can be defined as the convex hull of a finite number of points $P$ and is denoted by $\textrm{conv}(P)$, it is binary if all its extreme points are binary. A cone is a polyhedron described by a linear system with a right-hand side equal to 0; it is pointed if it contains no affine space. Given a set of points $P\subseteq\mathds{R}^{n}$, we denote by $\text{cone}(P)$ the conic hull of $P$’s elements, that is, the minimal cone containing $P$. A set of points $P$ is a generating set of the cone $\mathcal{C}$ if $\text{cone}(P)=\mathcal{C}$. The minimal generating sets of pointed cones are unique up to positive scalar multiplication; we call the elements of such a set its generators they are in one-to-one correspondence with its faces of dimension 1. A polyhedron $\mathcal{P}$ is said to have the integer decomposition property (IDP) if every integer point in the dilation $k\mathcal{P}$ for $k\in\mathds{Z}_{+}$ is obtained as the sum of $k$ elements of $\mathcal{P}$. A set of points $X=\{\chi_{1},\dots,\chi_{n}\}$ forms a Hilbert basis if every integer point in the conic hull of $X$ is expressible as an integer nonnegative combination of elements of $X$. Hilbert basis and integer decomposition property are closely related. If the extreme points of a binary polytope $\mathcal{P}$ form a Hilbert basis, then $\mathcal{P}$ has IDP.
@@
In this work, the polyhedra we consider are convex hulls of points encoding structured induced subgraphs of an input graph $G=(V,E)$. An induced subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ with $E^{\prime}=E(V^{\prime})$, can be encoded by its incidence vector $\chi^{V^{\prime}}\in\{0,1\}^{V}$ whose components associated with elements of $V^{\prime}$ are equal to 1 and the remaining components equal 0. In this article, we consider the extended incidence vector denoted by $\xi^{V^{\prime}}=(\chi^{V^{\prime}},\zeta^{E^{\prime}})\in\{0,1\}^{V}\times\{0,1\}^{E}$ where the components of $\zeta^{E^{\prime}}$ associated with elements of $E$ are equal to 1 and the remaining components equal 0. Considering extended descriptions such as proposed here is a common way to obtain compact polyhedral descriptions of combinatorial polytopes [exntededfo]. In particular, considering extended spaces sometimes allows to obtain compact descriptions, whereas the natural variable may require exponentially many constraints. Given a point $(x,y)\in\mathds{R}^{V}\times\mathds{R}^{E}$ we denote by $(x,y)|_{G^{\prime}}$ its restriction to elements of $G^{\prime}$.
Given a graph $G=(V,E)$, the polytopes we consider in this article are defined as follows:
  • its induced tree cone $\text{cone}\{\xi^{W}:\text{ $W$ is an induced tree of $G$}\}$
  • its induced tree polytope $\text{conv}\{\xi^{W}:\text{ $W$ is an induced tree of $G$}\}$
  • its induced path cone $\text{cone}\{\xi^{W}:\text{ $W$ is an induced path of $G$}\}$
  • its induced path polytope $\text{conv}\{\xi^{W}:\text{ $W$ is an induced path of $G$}\}$
Notations
Here and later, calligraphic letters denote polyhedra. Capital letters denote sets of vertices or edges, or graphs (note that $\delta$ is also used for sets of edges). Lowercase letters refer to vertices, edges or numbers. The symbol $\Gamma$ is reserved for sets of induced trees, while $\Psi$ for sets of induced paths, $\Omega$ for sets of cliques and $\lambda$ refers to vector of integer coefficients. Finally, $\Delta$ refers to gaps which we aim to reduce to 0.
@@
In this article, we study the problems of finding maximum weight induced trees and paths where both the vertices and the edges are weighted. In particular, we fill a gap in the polyhedral literature associated with these problems.
@@
The problem of finding the size of the largest induced tree in a graph is a problem that has been approached through extremal combinatorics [ERDOS198661, MATOUSEK2007307, FOX2009494]. In particular, an important question is that of bounding this number according to several graph parameters, such as the number of vertices, edges, radius, independence number, maximum clique, and connectivity. In [ERDOS198661], a Ramsey-type bound is proposed on the minimum integer $n(k,t)$ such that every graph having at least $n(k,t)$ vertices contains either a clique of size $k$ or an induced tree of size $t$. Finally, they show that every connected graph with radius $r$ has an induced tree of size at least 2r12r-1 by exhibiting a large induced path. A related problem is the one of finding large induced forests [inducedforest] in which bounds for the largest induced forest are given with respect to maximum degree or stability number. Besides these studies, no polyhedral study of the induced tree polytope on particular classes has been proposed. The problem of computing the maximum size of an induced tree is known to be NP-hard [completePathTree] in general, but it is also the case in bipartite and triangle-free graphs with maximal degree 3 [MATOUSEK2007307]. Several integer linear programs have been proposed for the maximum weight induced tree problem [ILPinducedtree] based on different variable spaces, and extensive computational experiments have been made on general graphs.
@@
The problem of finding a longest induced path has applications in assessing the robustness of a network. It can be viewed as a variant of the graph diameter. In a graph where vertices represent communication devices that may fail, and edges represent possible transmission links, the length of the longest induced path indicates the worst-case number of intermediate devices required for a one-to-one communication between any two vertices $u$ and $v$ after the failure of a set of devices that do not disconnect $u$ from $v$. Another application is about error correcting codes [error]. The induced paths of a graph are induced trees, and the type of questions raised in the previous paragraph can be raised considering this structure. The authors of [FOX2009494] show that every connected triangle-free graph on $n$ vertices has an induced path of length at least $\sqrt{n}$, and their approach can be generalized to obtain lower bounds for graphs with no clique of size $k$. In [esperet], the authors show that every 3-connected planar graph contains an induced path on $\text{log}(|V|)$ vertices. In [jfraymond], it is observed that a graph may contain a large path but no large induced path; for this reason, they characterize the sets of edges to be forbidden along a path of length $n$ to force the existence of an induced path of length $log(n)$. This result allows them to show that if a graph has a path of length $n$ and no $K_{t}$ as a topological minor, then it contains an induced path of order $log(n)^{\Omega(1/tlog^{2}t)}$.
Finding a maximum vertex weighted induced path is also NP-hard [completePathTree], but polynomially solvable in $k$-chordal graphs, interval filament, graphs decomposable by clique cutset or by splits whose prime subgraphs are polynomially solvable [GAVRIL2002203], improved in [HANSEN2009135] and on graphs of bounded clique width, asteroidal triple free graphs [inducedpathcliquewifth]. Besides these combinatorial algorithms, no polyhedral study of the corresponding polytope has been made. A related polyhedra has however, been considered in [goel2025halfspacerepresentationspathpolytopes]: the polytope whose extreme points in $\{0,1\}^{E}$ encode paths between pairs of leaves in a tree. In trees, induced path and path are the same subgraphs, so the latter article describes the induced path (between leaves) polytope of trees with edge variables. Finally, extended space integer linear programs have been proposed for the maximum weight induced path [ILPinducedPath, ILPinducedPath2], leading to computational experiments.
@@
In this article, we study both the induced tree polytope and the induced path polytopes of chordal graphs in two separate sections. It turns out that beside being included one in the other, these two polytopes share similar properties. Both following sections follow the same scheme, the first is about induced trees, and the second is about induced path. Each section starts by showing that the extended incidence vectors of their corresponding graph structure form a Hilbert basis in chordal graphs. As a byproduct, we obtain linear systems of inequalities describing the cones they generate in chordal graphs. Finally, we show that intersecting these cones with the correct hyperplane yields the induced tree and path polytope of chordal graphs in extended variable space. While the system we provide for the induced tree polytope is easily seen to be compact in chordal graphs, the one we provide for the induced path polytope has an exponential number of inequalities. Therefore, we characterize the ones that define facets and show that a superset can be enumerated in polynomial time. This implies that there exists a compact system describing the induced path polytope of chordal graphs. This study allows to solve the problems of finding a maximum vertex and edge weight induced tree or a maximum weight induced path in polynomial time for chordal graphs.